Given C= uso, a 10, and b 14, use the Jaw of Cosines to solve the triangle for the value of c.
Answer #1

Using Law of cosines c²- a²+6²_zab Cosc <?= (102+(1412-2(10) (14) (as 115 c? = 100+ 196 - 280 COS 1150 C = J 296 - 280 Cos 11